I am not sure if I am doing this problem the right way.
The formula for calculating the amount of money returned for an initial deposit money into a bank account or CD (Certificate of Deposit) is given by
A=P(1+r/n)^nt
A is the amount of returned.
P is the principal amount initially deposited.
r is the annual interest rate (expressed as a decimal).
n is the compound period.
t is the number of years.
Suppose you deposit $10,000 for 2 years at a rate of 10%.
Calculate the return (A) if the bank compounds quarterly (n = 4). Round your answer to the hundredth's place.
Thanks for your time!

Answer by rapaljer(4671)

If the bank compounds quarterly, that is 4 times per year, so n = 4, t=2, and nt = 8.
A=P(1+r/n)^nt
A=10000(1+.10/4)^8
A=10000(1.025)^8
A=10000(1.2184029)
A = $12184.03

R^2 at SCC